OpenStudy (anonymous):

can someone help me plz :) The midpoint of a line segment with end points as (x, 4) and (-2, -2) is (2, 1). What is the value of x? x = 6 x = 4 x = -3 x = -2

OpenStudy (anonymous):



OpenStudy (anonymous):

Midpoint = total/2 2 = ( x -2)/ 2 @nataliem Either way will work :)

OpenStudy (anonymous):

cool thanx alot :)

OpenStudy (anonymous):

@nataliem What's your result?

OpenStudy (anonymous):

still trying to solve

OpenStudy (anonymous):

i can solve it i keep getting a weird ans

OpenStudy (anonymous):

* cant

OpenStudy (anonymous):

I'd like to see your mistake so I can fix it :)

OpenStudy (anonymous):

okay 2,1=x=-2/2 , 4=-2/2 thats the formula right?

OpenStudy (anonymous):

2,1=x+-2/2, 4+-2/2 sorry bout above

OpenStudy (anonymous):

Have you read my post ???

OpenStudy (anonymous):

u men this ? Midpoint = total/2 2 = ( x -2)/ 2 @nataliem Either way will work :)

OpenStudy (anonymous):

Yes, now solve for x!

OpenStudy (anonymous):

its 6

OpenStudy (anonymous):

Finally, after 1 hour! Yes, x = 6